# Optimal Rates in Continual Linear Regression via Increasing Regularization

### v1

@@ Appendix A Proofs of lower bounds
Let $d\geq 2$ and $k\geq 2$. Then for any algorithm $\mathcal{A}$ which receives $k$ functions $f_{1},f_{2},\ldots,f_{k}:\mathbb{R}^{d}\to\mathbb{R}$ and outputs a point in $\mathbb{R}^{d}$, there exists a point ${\mathbf{w}}_{\star}\in\mathbb{R}^{d}$ such that $\left\|{{\mathbf{w}}_{\star}}\right\|\leq 1$ and a set of $k$ $1$-smooth convex quadratic functions which are minimized at ${\mathbf{w}}_{\star}$, $h_{1},\ldots,h_{k}:\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Lemma A.2 (Sherman-Morison).
@@
  • Recall Lemma 4.9 — lower bound for seen-task loss under Scheme 1.For any $d\geq 2$, initialization ${\mathbf{w}}_{0}\in\mathbb{R}^{d}$, and regularization coefficient sequence $\lambda_{1},\ldots,\lambda_{k}\geq 0$, there exists a set of jointly realizable linear regression tasks $\left\{\left({\mathbf{X}_{m},\mathbf{y}_{m}}\right)\right\}_{m=1}^{M}$ such that, under a with-replacement random task ordering, Scheme 1 incurs seen-task loss $\mathcal{L}_{1:k}({\mathbf{w}}_{k})=\Omega(1/k)$ with probability at least $1/10$.
@@
    where the second equality follows from Lemma A.2. The case of $\lambda=0$ is treated as the update above with $\lambda=0$, and similarly,
@@
    and $f_{1}({\mathbf{w}}_{k})=\frac{1}{2}\zeta^{2}.$ If $|\zeta|\geq\frac{1}{\sqrt{k}}$, we are done as $f_{1}$ is observed $k-1$ times (conditioned on $E_{B}$) and
@@
Appendix B Proofs of the reductions and their properties
@@
Lemma B.1 (General reduction properties).
@@
where $g:\mathbb{R}\to\mathbb{R}$ is applied spectrally (i.e., to each eigenvalue of $\mathbf{X}_{m}^{\top}\mathbf{X}_{m}$). Assume that $g$ is concave, non-decreasing on $[0,R^{2}]$, with $g(0)=0$ and $g^{\prime}(0)>0$. Then:
@@
Hence, $\frac{1}{g^{\prime}(0)}\mathbf{A}_{m}\preccurlyeq\mathbf{X}_{m}^{\top}\mathbf{X}_{m}.$ By concavity and $g(0)=0$, the chord from 00 to $R^{2}$ lies below $g$:
@@
  • Recall Lemma 3.1 — properties of the IGD objectives.For $t\in\left[{k}\right]$, define $f^{(t)}_{r},f^{(t)}_{b}$ as in Reductions 1 and 2, and recall the data radius $R\triangleq\max_{m\in\left[{M}\right]}\left\|{\mathbf{X}_{m}}\right\|_{2}$.
@@
      Finally, when the tasks are jointly realizable (see Assumption 4.1), the same ${\mathbf{w}}_{\star}$ minimizes all surrogate objectives simultaneously. That is,
@@
    We verify that $g_{r}$ and $g_{b}$ satisfy the assumptions of Lemma B.1 on the domain $\xi\in[0,R^{2}]$:
@@
    Hence, by Lemma B.1, both $f^{(t)}_{r}$ and $f^{(t)}_{b}$ are $\beta^{(t)}$-smooth with the claimed parameters $\beta_{r}^{(t)},\beta_{b}^{(t)}$, and they satisfy the two-sided bounds:
@@
    Finally, for part (iii), assume the tasks satisfy joint realizability (Assumption 4.1), meaning that for some common minimizer ${\mathbf{w}}_{\star}$,
@@
Appendix C Proofs for fixed regularization strength
  • Recall Lemma 4.3 — rates for fixed regularization strength.Assume a random with-replacement ordering over jointly realizable tasks. Then, for each of Schemes 1 and 2, the expected loss after $k\geq 1$ iterations is upper bounded as:
@@ Appendix C Proofs for fixed regularization strength
      Fixed coefficient: For Scheme 1 with a regularization coefficient $\lambda>0$,
@@
      Fixed budget: For Scheme 2 with step size $\gamma\in(0,1/R^{2})$ and budget $N\in\mathbb{N}$,
@@
    Proof of Lemma 4.3.
    From Reductions 1 and 2, the iterates of Schemes 1 and 2 are equivalent to those of IGD (Scheme 3) applied to the respective surrogate objectives $f^{(t)}_{r}$ and $f^{(t)}_{b}$. When $\eta,\lambda,\gamma,N$ are fixed, the functions $f^{(t)}_{r},f^{(t)}_{b}$ do not depend on $t$, and under a random ordering with replacement, the update rule becomes standard SGD.
    By Lemma 3.1, the surrogates $f^{(t)}_{r}$ and $f^{(t)}_{b}$ are jointly realizable whenever the original losses are, and hence satisfy the assumptions of the following result from Evron et al. [13].
@@
    (i) Fixed Regularization. For Scheme 1, the surrogate $f^{(t)}_{r}$ is $\beta_{r}$-smooth with
@@
    (ii) Fixed Budget. For Scheme 2, the surrogate $f^{(t)}_{b}$ is $\beta_{b}$-smooth with
@@
    To extend this result to the without-replacement case (see Remark 4.5), we can simply invoke the without-replacement extension of Theorem 5.1 in Evron et al. [13]. ∎
  • Recall Corollary 4.4 — near-optimal rates via fixed regularization strength.Assume a random with-replacement ordering over jointly realizable tasks. When the regularization strengths in Lemma 4.3 are set as follows:
@@
      Fixed coefficient: For Scheme 1, set regularization coefficient $\lambda\triangleq R^{2}(\ln k-1)$;
@@
      Fixed budget: For Scheme 2, choose step size $\gamma\in(0,1/R^{2})$ and set budget $N\triangleq\frac{\ln\left({1-\frac{1}{\ln k}}\right)}{\ln\left({1-\gamma R^{2}}\right)}$;
    Then, under either Scheme 1 or Scheme 2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
    Proof of Corollary 4.4.
    We apply the general loss bound from Lemma 4.3, which holds for both fixed-regularization and fixed-budget variants:
@@
Appendix D Proofs for scheduled regularization strength
  • Recall Theorem 4.6 — optimal rates for increasing regularization.Assume a random with-replacement ordering over jointly realizable tasks. Consider either Scheme 1 or Scheme 2 with the following time-dependent schedules:
@@ Appendix D Proofs for scheduled regularization strength
      Scheduled coefficient: For Scheme 1, set regularization coefficient $\lambda_{t}=\frac{13R^{2}}{3}\cdot\frac{k+1}{k-t+2}$;
@@
      For Scheme 2, choose step sizes $\gamma_{t}\in(0,1/R^{2})$ and set budget $N_{t}=\frac{3}{13\gamma_{t}R^{2}}\cdot\frac{k-t+2}{k+1}$;
    Then, under either Scheme 1 or Scheme 2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
    Proof of Theorem 4.6.
    We apply Lemma 4.7 with the original loss $f({\mathbf{w}};m)=\mathcal{L}({\mathbf{w}};m)$ and surrogates $f^{(t)}({\mathbf{w}};m)=f^{(t)}_{r}({\mathbf{w}};m)$ or $f^{(t)}_{b}({\mathbf{w}};m)$, defined in Reductions 1 and 2.
    Smoothness and convexity. From Lemma 3.1, both surrogates are convex. Their smoothness constants are:
@@
    Thus, both surrogates are $R^{2}$-smooth, matching the smoothness of the loss $\mathcal{L}(\cdot;m)$ and satisfying condition (i) of Lemma 4.7.
    Joint realizability. From Lemma 3.1, if the original tasks are jointly realizable, then so are the surrogates:
@@
    so condition (iii) of Lemma 4.7 is satisfied.
    Two-sided bounds. We verify condition (ii) of Lemma 4.7 using the two-sided inequalities from Lemma 3.1:
@@
    Applying Lemma 4.7 yields:
@@
D.1 Proof of Lemma 4.7
In this section, we provide the proof of our main lemma establishing the guarantees of time varying SGD. In order to better align with conventions in the optimization literature from which our techniques draw upon, we adopt different indexing for the SGD iterates throughout this section. Below, we restate the lemma with the alternative indexing scheme; the original Lemma 4.7 follows immediately by a simple shift of $k+1\to k$ and $1\to 0$ in the indexes of the iterates ${\mathbf{w}}_{t}$.
Lemma D.1 (Restatement of Lemma 4.7 with alternative indexing).
Assume $\tau$ is a random with-replacement ordering over $M$ jointly-realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. Define the average loss $f({\mathbf{w}})\triangleq\mathbb{E}_{m\sim\tau}f({\mathbf{w}};m)$. Let $k\geq 2$, and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[k],m\in[M]\right\}$ for $t\in[k]$ are time-varying surrogate losses that satisfy:
@@
Theorem D.2.
Let $J\geq 2$, and assume $\tau\colon[J]\to[M]$ is a random with-replacement ordering over $M$ jointly-realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[J],m\in[M]\right\}$ for $t\in[J]$ are time-varying surrogate losses for which there exists a weight sequence $\nu_{1},\ldots,\nu_{J}$ that satisfies, for all $m\in[M],t\in[J],{\mathbf{w}}\in\mathbb{R}^{d}$:
@@
Next, we prove a technical lemma which we employ in conjunction with the above in the proof of Lemma D.1.
Lemma D.3.
@@
The next lemma provides the stability property, which we leverage to translate our loss guarantees to the seen-task loss defined in Definition 4.8.
Lemma D.4.
Assume the conditions of Lemma D.1 and consider the algorithm defined in Eq. 1 with non-increasing step sizes $\eta_{t}\leq 1/2\beta$. In addition, define for every $1\leq k$, $\hat{f}_{1:k}({\mathbf{w}})\triangleq\frac{1}{k}\sum_{t=1}^{k}f({\mathbf{w}};\tau_{t})$. For all $1\leq k$, the following holds:
@@
Next, we bound $\left\|{{\mathbf{w}}_{k+1}^{(i)}-{\mathbf{w}}_{k+1}}\right\|^{2}$. Since by Lemma D.1, for every $t$, $f^{(t)}$ is convex and $\beta$-smooth, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when for every $t$, $\eta_{t}\leq 2/\beta$ [see Lemma 3.6 in 18]:
@@
In addition, denoting by $f_{m^{\prime}_{i}}$ the function that sampled after replacing $f_{{m}_{i}}$ and its corresponding time varying objective by $f^{(m^{\prime}_{i})}$, by the conditions in Lemma D.1, we have that,
@@
Now, by Theorem D.2 with $v_{t}=1$ for every $t$, we have, since $\eta_{t}\beta\leq\frac{1}{4}$, $\frac{1}{1+\eta_{t}\beta}\geq\frac{4}{5}$
@@
Proof of Lemma D.1.
To begin, note that we are after a guarantee for ${\mathbf{w}}_{k+1}$, which is the SGD iterate that was produced by taking $k$ steps over $k$ losses. To that end, we are going to apply Theorem D.2 with $J=k+1$, hence we are obligated to supply a random ordering $\tau\colon[k+1]\to[M]$, $f^{(k+1)}$ and $\eta_{k+1}$, which are not supplied in the statement of our lemma. Therefore, we define
@@
Now, the conditions for Theorem D.2 are immediately satisfied with $J=k+1$ by our assumptions and augmentation described above, leading to:
@@
Now, by Lemma D.3, the step size sequence $\eta_{t}=\eta(\frac{(k+1)-t+1}{k+1})$ with $\eta\leq\frac{3}{13\beta}$ and $\left\{v_{t}\right\}_{t=1}^{k+1}$ as specified by the lemma, guarantee that $c_{k+1}\geq\frac{\eta}{k+1}$, $v_{0}\leq 3/(k+1)$, and $c_{t}\geq 0$ for all $t\in[k+1]$. Combining these properties with Section D.1 we obtain,
@@
which completes the proof for the first part. For the seen-task guarantee, by Lemma D.4, we have

### v2

@@ Appendix B Proofs of lower bounds
Let $d\geq 2$ and $k\geq 2$. Then for any algorithm $\mathcal{A}$ which receives $k$ functions $f_{1},f_{2},\ldots,f_{k}:\mathbb{R}^{d}\to\mathbb{R}$ and outputs a point in $\mathbb{R}^{d}$, there exists a point ${\mathbf{w}}_{\star}\in\mathbb{R}^{d}$ such that $\left\|{{\mathbf{w}}_{\star}}\right\|\leq 1$ and a set of $k$ 11-smooth convex quadratic functions which are minimized at $\mathbf{w}$$_{\star}$, $h_{1},\ldots,h_{k}:\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Lemma B.2 (Sherman-Morison).
@@
Recall Lemma˜4.9 — lower bound for seen-task loss under Scheme˜1.For any $d\geq 2$, initialization ${\mathbf{w}}_{0}\in\mathbb{R}^{d}$, and regularization coefficient sequence $\lambda_{1},\ldots,\lambda_{k}\geq 0$, there exists a set of jointly realizable linear regression tasks $\left\{\left({\mathbf{X}_{m},\mathbf{y}_{m}}\right)\right\}_{m=1}^{M}$ such that, under a with-replacement random task ordering, Scheme˜1 incurs seen-task loss $\mathcal{L}_{1:k}({\mathbf{w}}_{k})=\Omega(1/k)$ with probability at least $1/10$.
@@
where the second equality follows from Lemma˜B.2. The case of $\lambda=0$ is treated as the update above with $\lambda=0$, and similarly,
@@
and $f_{1}({\mathbf{w}}_{k})=\frac{1}{2}\zeta^{2}.$ If $|\zeta|\geq\frac{1}{\sqrt{k}}$, we are done as $f_{1}$ is observed k1k-1 times (conditioned on $E_{B}$) and
@@
Appendix C Proofs of the reductions and their properties
@@
Lemma C.1 (General reduction properties).
@@
where $g:\mathbb{R}\to\mathbb{R}$ is applied spectrally (i.e., to each eigenvalue of $\mathbf{X}_{m}^{\top}\mathbf{X}_{m}$). Assume that $g$ is concave, non-decreasing on $0,R^{2}]$, with $g(0)=0$ and $g^{\prime}(0)>0$. Then:
@@
Hence, $\frac{1}{g^{\prime}(0)}\mathbf{A}_{m}\preccurlyeq\mathbf{X}_{m}^{\top}\mathbf{X}_{m}.$ By concavity and $g(0)=0$, the chord from 0 to $R^{2}$ lies below $g$:
@@
Recall Lemma˜3.1 — properties of the IGD objectives.For $t\in\left[{k}\right]$, define $f^{(t)}_{r},f^{(t)}_{b}$ as in Reductions˜1 and 2, and recall the data radius $R\triangleq\max_{m\in\left[{M}\right]}\left\|{\mathbf{X}_{m}}\right\|_{2}$.
@@
    Finally, when the tasks are jointly realizable (see Assumption˜4.1), the same $\mathbf{w}$$_{\star}$ minimizes all surrogate objectives simultaneously. That is,
@@
We verify that $g_{r}$ and $g_{b}$ satisfy the assumptions of Lemma˜C.1 on the domain $\xi\in[0,R^{2}]$:
@@
Hence, by Lemma˜C.1, both $f^{(t)}_{r}$ and $f^{(t)}_{b}$ are $\beta^{(t)}$-smooth with the claimed parameters $\beta_{r}^{(t)},\beta_{b}^{(t)}$, and they satisfy the two-sided bounds:
@@
Finally, for part (iii), assume the tasks satisfy joint realizability (Assumption˜4.1), meaning that for some common minimizer $\mathbf{w}$$_{\star}$,
@@
Appendix D Proofs for fixed regularization strength
Recall Lemma˜4.3 — rates for fixed regularization strength.Assume a random with-replacement ordering over jointly realizable tasks. Then, for each of Schemes˜1 and 2, the expected loss after $k\geq 1$ iterations is upper bounded as:
@@ Appendix D Proofs for fixed regularization strength
    Fixed coefficient: For Scheme˜1 with a regularization coefficient $\lambda>0$,
@@
    Fixed budget: For Scheme˜2 with step size $\gamma\in(0,1/R^{2})$ and budget $N\in\mathbb{N}$,
@@
Proof of Lemma˜4.3.
From Reductions˜1 and 2, the iterates of Schemes˜1 and 2 are equivalent to those of IGD (Scheme˜3) applied to the respective surrogate objectives $f^{(t)}_{r}$ and $f^{(t)}_{b}$. When $\eta,\lambda,\gamma,N$ are fixed, the functions $f^{(t)}_{r},f^{(t)}_{b}$ do not depend on $t$, and under a random ordering with replacement, the update rule becomes standard SGD.
By Lemma˜3.1, the surrogates $f^{(t)}_{r}$ and $f^{(t)}_{b}$ are jointly realizable whenever the original losses are, and hence satisfy the assumptions of the following result from Evron et al. [14].
@@
(i) Fixed Regularization. For Scheme˜1, the surrogate $f^{(t)}_{r}$ is $\beta_{r}$-smooth with
@@
(ii) Fixed Budget. For Scheme˜2, the surrogate $f^{(t)}_{b}$ is $\beta_{b}$-smooth with
@@
To extend this result to the without-replacement case (see Remark˜4.5), we can simply invoke the without-replacement extension of Theorem 5.1 in Evron et al. [14]. ∎
Recall Corollary˜4.4 — near-optimal rates via fixed regularization strength.Assume a random with-replacement ordering over jointly realizable tasks. When the regularization strengths in Lemma˜4.3 are set as follows:
@@
    Fixed coefficient: For Scheme˜1, set regularization coefficient $\lambda\triangleq R^{2}(\ln k-1)$;
@@
    Fixed budget: For Scheme˜2, choose step size $\gamma\in(0,1/R^{2})$ and set budget $N\triangleq\frac{\ln\left({1-\frac{1}{\ln k}}\right)}{\ln\left({1-\gamma R^{2}}\right)}$;
Then, under either Scheme˜1 or Scheme˜2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
Proof of Corollary˜4.4.
We apply the general loss bound from Lemma˜4.3, which holds for both fixed-regularization and fixed-budget variants:
@@
Appendix E Proofs for scheduled regularization strength
Recall Theorem˜4.6 — optimal rates for increasing regularization.Assume a random with-replacement ordering over jointly realizable tasks. Consider either Scheme˜1 or Scheme˜2 with the following time-dependent schedules:
@@ Appendix E Proofs for scheduled regularization strength
    Scheduled coefficient: For Scheme˜1, set regularization coefficient $\lambda_{t}=\frac{13R^{2}}{3}\cdot\frac{k+1}{k-t+2}$;
@@
    For Scheme˜2, choose step sizes $\gamma_{t}\in(0,1/R^{2})$ and set budget $N_{t}=\frac{3}{13\gamma_{t}R^{2}}\cdot\frac{k-t+2}{k+1}$;
Then, under either Scheme˜1 or Scheme˜2, the expected loss after $k\geq 2$ iterations is bounded as:
@@
Proof of Theorem˜4.6.
We apply Lemma˜4.7 with the original loss $f({\mathbf{w}};m)=\mathcal{L}({\mathbf{w}};m)$ and surrogates $f^{(t)}({\mathbf{w}};m)=f^{(t)}_{r}({\mathbf{w}};m)$ or $f^{(t)}_{b}({\mathbf{w}};m)$, defined in Reductions˜1 and 2.
Smoothness and convexity. From Lemma˜3.1, both surrogates are convex. Their smoothness constants are:
@@
Thus, both surrogates are $R^{2}$-smooth, matching the smoothness of the loss $\mathcal{L}(\cdot;m)$ and satisfying condition (i) of Lemma˜4.7.
Joint realizability. From Lemma˜3.1, if the original tasks are jointly realizable, then so are the surrogates:
@@
so condition (iii) of Lemma˜4.7 is satisfied.
Two-sided bounds. We verify condition (ii) of Lemma˜4.7 using the two-sided inequalities from Lemma˜3.1:
@@
Applying Lemma˜4.7 yields:
@@
E.1 Proof of Lemma˜4.7
In this section, we provide the proof of our main lemma establishing the guarantees of time varying SGD. In order to better align with conventions in the optimization literature from which our techniques draw upon, we adopt different indexing for the SGD iterates throughout this section. Below, we restate the lemma with the alternative indexing scheme; the original Lemma˜4.7 follows immediately by a simple shift of $k+1\to k$ and $1\to 0$ in the indexes of the iterates ${\mathbf{w}}_{t}$.
Lemma E.1 (Restatement of Lemma˜4.7 with alternative indexing).
Assume $\tau$ is a random with-replacement ordering over $M$ jointly realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. Define the average loss $f({\mathbf{w}})\triangleq\mathbb{E}_{m\sim\tau}f({\mathbf{w}};m)$. Let $k\geq 2$, and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[k],m\in[M]\right\}$ for $t\in[k]$ are time-varying surrogate losses that satisfy:
@@
Theorem E.2.
Let $J\geq 2$, and assume $\tau\colon[J]\to[M]$ is a random with-replacement ordering over $M$ jointly realizable convex and $\beta$-smooth loss functions $f(\cdot;m)\colon\mathbb{R}^{d}\to\mathbb{R}$. and suppose $\left\{f^{(t)}(\cdot;m)\mid t\in[J],m\in[M]\right\}$ for $t\in[J]$ are time-varying surrogate losses for which there exists a weight sequence $\nu_{1},\ldots,\nu_{J}$ that satisfies, for all $m\in[M],t\in[J],{\mathbf{w}}\in\mathbb{R}^{d}$:
@@
Next, we prove a technical lemma which we employ in conjunction with the above in the proof of Lemma˜E.1.
Lemma E.3.
@@
The next lemma provides the stability property, which we leverage to translate our loss guarantees to the seen-task loss defined in Definition˜4.8.
Lemma E.4.
Assume the conditions of Lemma˜E.1 and consider the algorithm defined in Eq.˜1 with non-increasing step sizes $\eta_{t}\leq 1/2\beta$. In addition, define for every $1\leq k$, $\hat{f}_{1:k}({\mathbf{w}})\triangleq\frac{1}{k}\sum_{t=1}^{k}f({\mathbf{w}};\tau_{t})$. For all $1\leq k$, the following holds:
@@
Next, we bound $\left\|{{\mathbf{w}}_{k+1}^{(i)}-{\mathbf{w}}_{k+1}}\right\|^{2}$. Since by Lemma˜E.1, for every $t$, $f^{(t)}$ is convex and $\beta$-smooth, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when for every $t$, $\eta_{t}\leq 2/\beta$ [see Lemma 3.6 in 19]:
@@
In addition, denoting by $f_{m^{\prime}_{i}}$ the function that sampled after replacing $f_{{m}_{i}}$ and its corresponding time varying objective by $f^{(m^{\prime}_{i})}$, by the conditions in Lemma˜E.1, we have that,
@@
Now, by Theorem˜E.2 with $v_{t}=1$ for every $t$, we have, since $\eta_{t}\beta\leq\frac{1}{4}$, $\frac{1}{1+\eta_{t}\beta}\geq\frac{4}{5}$
@@
Proof of Lemma˜E.1.
To begin, note that we are after a guarantee for ${\mathbf{w}}_{k+1}$, which is the SGD iterate that was produced by taking $k$ steps over $k$ losses. To that end, we are going to apply Theorem˜E.2 with $J=k+1$, hence we are obligated to supply a random ordering $\tau\colon[k+1]\to[M]$, $f^{(k+1)}$ and $\eta_{k+1}$, which are not supplied in the statement of our lemma. Therefore, we define
@@
Now, the conditions for Theorem˜E.2 are immediately satisfied with $J=k+1$ by our assumptions and augmentation described above, leading to:
@@
Now, by Lemma˜E.3, the step size sequence $\eta_{t}=\eta(\frac{(k+1)-t+1}{k+1})$ with $\eta\leq\frac{3}{13\beta}$ and $\left\{v_{t}\right\}_{t=1}^{k+1}$ as specified by the lemma, guarantee that $c_{k+1}\geq\frac{\eta}{k+1}$, $v_{0}\leq 3/(k+1)$, and $c_{t}\geq 0$ for all $t\in[k+1]$. Combining these properties with Section˜E.1 we obtain,
@@
which completes the proof for the first part. For the seen-task guarantee, by Lemma˜E.4, we have